OpenStudy (anonymous):

Simplify: 1 -------- -27^-2/3

OpenStudy (anonymous):

use properties: \[x ^{-b}= 1/x ^{b}\] and \[x ^{c/d}= [\sqrt[d]{x}]^{c}\]